# Cross-Problem Learning for Solving Vehicle Routing Problems

### v1

@@ Abstract
Existing neural heuristics often train a deep architecture from scratch for each specific vehicle routing problem (VRP), ignoring the transferable knowledge across different VRP variants. This paper proposes the cross-problem learning to assist heuristics training for different downstream VRP variants. Particularly, we modularize neural architectures for complex VRPs into 1) the backbone Transformer for tackling the travelling salesman problem (TSP), and 2) the additional lightweight modules for processing problem-specific features in complex VRPs. Accordingly, we propose to pre-train the backbone Transformer for TSP, and then apply it in the process of fine-tuning the Transformer models for each target VRP variant. On the one hand, we fully fine-tune the trained backbone Transformer and problem-specific modules simultaneously. On the other hand, we only fine-tune small adapter networks along with the modules, keeping the backbone Transformer still. Extensive experiments on typical VRPs substantiate that 1) the full fine-tuning achieves significantly better performance than the one trained from scratch, and 2) the adapter-based fine-tuning also delivers comparable performance while being notably parameter-efficient. Furthermore, we empirically demonstrate the favorable effect of our method in terms of cross-distribution application and versatility.
@@ 1 Introduction
  • As an early attempt, we propose cross-problem learning for solving VRPs. The pre-trained backbone Transformer for a basic VRP (i.e., TSP) is used to foster the training of Transformers for downstream complex VRPs through fine-tuning.
@@
  • We develop different fine-tuning methods for cross-problem learning, i.e., the full and adapter-based fine-tuning, by which we fine-tune the entire Transformers or only the lightweight problem-specific modules along with small adapter networks.
  • We empirically testify that the knowledge learned in the Transformer for TSP is well transferred to aid in training neural heuristics for other VRPs. While full fine-tuning achieves better performance than Transformer trained from scratch, adapter-based fine-tuning methods attain comparable performance, with far fewer parameters to be trained and stored.
@@ 2.1 Neural Heuristics for VRP
Current neural heuristics for VRPs mostly learn policies to construct a solution in an autoregressive way. Among diverse neural construction heuristics Bello et al. (2017); Nazari et al. (2018); Dai et al. (2017), the first breakthrough is made by the Transformer Vaswani et al. (2017), which is cast as Attention Model (AM) to solve VRP variants such as TSP, OP and PCTSP Kool et al. (2018). After that, numerous neural heuristics are proposed based on AM for solving VRPs Xin et al. (2021a); Kwon et al. (2020); Hottung et al. (2022). Among them, Kwon et al. significantly ameliorate AM with policy optimization with multiple optima (POMO) Kwon et al. (2020), and enlighten considerable follow-up works with Transformers to solve VRPs Kim et al. (2022b); Kwon et al. (2021); Choo et al. (2022). As alternatives, some works endeavor to predict probabilities of edges in optimal solutions to VRPs, which are then leveraged to efficiently construct high-quality solutions with search algorithms such as greedy search, sampling, and Monte Carlo tree search Fu et al. (2021); Qiu et al. (2022); Hudson et al. (2022). On the other hand, neural improvement heuristics aim to learn policies to iteratively improve an initial but complete solution. Inspired by traditional local search algorithms, a few works learn local operations such as 2-opt and swap to promote the solution  Wu et al. (2021); Ma et al. (2021); Wang et al. (2021); Kim et al. (2022a). Inspired by specialized algorithms or solvers for VRPs, some literature attempts to learn repair or destroy operations in Large Neighborhood Search (LNS) Hottung and Tierney (2020); Gao et al. (2020). Deep learning is also applied to enhance solvers like Lin-Kernighan-Helsgaun (LKH) Xin et al. (2021b); Kim et al. (2021) and Hybrid Genetic Search (HGS) Santana et al. (2022).
@@ 2.2 Pre-training, Fine-tuning
The pre-training-then-fine-tuning paradigm has been widely verified effective in domains of Natural Language Processing (NLP) Lewis et al. (2020); Dong et al. (2019); Devlin et al. (2018), Graph Neural Network (GNN) Qiu et al. (2020); Yun et al. (2019); Hu et al. (2020), Computer Vision (CV) Bao et al. (2022); He et al. (2020); Chen et al. (2020), etc. In this line of work, large models such as Transformers are first trained on basic tasks Dong et al. (2019); Ribeiro et al. (2020); Hu et al. (2020); You et al. (2020); Dosovitskiy et al. (2021); Touvron et al. (2021). Then the pre-trained model is fine-tuned to solve downstream tasks, with the neural network adapted by introducing additional structures or layers (which possess far fewer parameters than the pre-trained model) Yun et al. (2019); Houlsby et al. (2019); Pfeiffer et al. (2021); Zhao et al. (2020); Kitaev et al. (2020); Hu et al. (2021). In this manner, the pre-training can be conducted on considerable unlabeled data to learn general patterns and features, which are re-used in downstream tasks. Also, the knowledge learned during pre-training is transferred to fine-tuning process to facilitate more efficient training.
@@ 3.1 VRP Description
Feasibility. The solution to a VRP instance is feasible if it satisfies the problem-specific constraints. For example, a feasible solution to a TSP instance is a tour that goes through every node on the graph exactly once. A solution to OP is feasible if the length of a tour is not larger than a given value. Other VRPs like PCTSP and CVRP are defined on graphs in a similar fashion. Their details can be found in Appendix A.
@@ 3.2 Transformer based Construction Heuristics
$\footnotesize\nabla_{\theta}\mathcal{L}(\theta|\mathcal{G})=\mathbb{E}_{p_{\theta}(\pi|\mathcal{G})}[(\mu(\pi)-b(\mathcal{G}))\nabla_{\theta}\log p_{\theta}(\pi|\mathcal{G})],$ (1)
@@ 4.4 Adapter-based Fine-tuning for VRPs
Inside Tuning. The adapter network in the inside tuning comprises two linear layers with a LeakyReLU (LR) activation function in between such that:
$\footnotesize\text{APT}_{\text{in}}(h_{i})=\mathbf{W}_{1}^{\text{in}}\cdot\text{LR}(\mathbf{W}_{0}^{\text{in}}\cdot h_{i}+\mathbf{b}_{0}^{\text{in}})+\mathbf{b}_{1}^{\text{in}},$ (2)
@@
Side Tuning. Instead of adjusting intermediate embeddings in the encoder, the side tuning adjusts the output from the encoder, with the following neural architecture:
$\footnotesize h^{{}^{\prime}}_{i}=\mathbf{BN}\left(h_{i}+\mathbf{MSL}_{i}\left(h_{0},\ldots,h_{n}\right)\right),$ (3)
$\footnotesize h_{i}^{{}^{\prime\prime}}=\mathbf{BN}\left(h^{{}^{\prime}}_{i}+\mathbf{FL}(h^{{}^{\prime}}_{i})\right),$ (4)
$\footnotesize\text{APT}_{\text{si}}(h_{i}^{{}^{\prime\prime}})=\mathbf{BN}(\text{LR}(\mathbf{W}_{1}^{\text{si}}\cdot\mathbf{BN}(\text{LR}(\mathbf{W}_{0}^{\text{si}}\cdot h_{i}^{{}^{\prime\prime}}+\mathbf{b}_{0}^{\text{si}}))+\mathbf{b}_{1}^{\text{si}})),$ (5)
@@
LoRA. Inspired by Low-Rank Adaptation (LoRA) in Hu et al. (2021), we design a low-rank decomposition to adjust the output from any matrix in the pre-trained Transformer,
$\footnotesize\text{APT}_{\text{lo}}(h_{i})=\mathbf{W}_{p}\cdot h_{i}+\mathbf{B}_{p}\mathbf{A}_{p}\cdot h_{i},$ (6)
@@ 5 Experiments
Settings. Following the existing works, we generate the coordinates of nodes by uniformly sampling from the unit square $[0,1]\times[0,1]$. For OP, we set prizes all equal to 1. For PCTSP, we randomly sample prizes and penalties from Uniform distributions, following Kool et al. (2018). Meanwhile, we generate different-sized instances for OP and PCTSP, by setting the number of customer nodes to $n=$ 20, 50, and 100. We then use the Transformer based AM in Kool et al. (2018) to implement the cross-problem learning, and maintain most of the original experimental settings for a fair comparison. Specifically, we train the model with 2500 batches in each epoch and use 512 instances in each batch (except for the problem with $n=$ 100, where we use 256 instances). For each epoch, the training instances are generated on the fly, and we use 10,000 instances for the validation. Then the paired t-test ($\alpha=5\%$) is conducted after each epoch to replace the baseline model, according to the improvement significance. For all problems and sizes, we train 100 epochs for the full convergence of the training curves. The learning rate is set as $10^{-4}$ in all our methods, as the one for the original AM [^1] We tried to assign different learning rates to adapter networks and problem-specific modules for fine-tuning, which did not result in better performance.. The server with Intel Core i9-10940X CPU and NVIDIA GeForce RTX 2080Ti is used in our experiments.
Baselines. We compare our method with AM Kool et al. (2018), the DRL based method with a strong performance on a variety of VRPs. We deploy the cross-problem learning based on its Transformer models. We also compare various baselines as follows. Regarding OP, we include 1) Gurobi, the state-of-the-art exact solver for various COPs Gurobi Optimization, LLC (2022); 2) Compass, the specialized Genetic Algorithm (GA) for OP Kobeaga et al. (2018); 3) Tsili, a classic randomized construction heuristic for OP with the manually crafted node probabilities  Tsiligirides (1984). As for PCTSP, besides Gurobi, we run two additional baselines: 4) ILS, the iterated local search metaheuristic that is widely used for routing problems; and 5) OR-Tools, a mature metaheuristics based solver for routing problems.
Table 1: Comparison between cross-problem learning and baselines.
@@
    Bold means the best results among learning based methods. * means the results of the method are from Kool et al. (2018).
Figure 3: Ablation study on fine-tuning methods and activation functions.
@@ 5.1 Comparison
We test all methods with 10,000 instances of OP and PCTSP, two typical VRPs that aim to maximize the total prize and minimize the distance plus the total penalty in the tour, respectively. We report the results of AM and Tsili by two decoding strategies, i.e., greedy search and sampling. For our method, we only report the better results derived by sampling. The sampling size is 1,280 for these methods, following Kool et al. (2018). For adapter-based fine-tuning, we only report results derived by inside tuning, since it significantly outstrips side tuning and LoRA (see Section 5.2). We compare the methods by different metrics, including the average objective value, average gap (to the best solutions), total runtime, and the number of trained parameters (for learning based methods). In Table 1, we can observe that our method with either full fine-tuning or inside tuning consistently outperforms AM with either greedy search or sampling, with significantly smaller gaps and similar runtime. While Gurobi achieves the optimal solutions for small problems, it cannot derive solutions for middle-sized OP and PCTSP ($n$=50, 100). Instead, we run Gurobi as a heuristic with a 30s limit, but it is inferior to our methods on OP ($n$=50, 100) or fails to obtain any solutions on PCTSP ($n$=100). While Compass and ILS are specialized algorithms for OP and PCTSP respectively, our method achieves comparable results in all cases, with most gap differences smaller than $1\%$. Moreover, our methods significantly surpass Tsili on OP ($n$=20, 50, 100), and outperform OR-Tools on PCTSP ($n$=100). Lastly, the inside tuning costs far fewer parameters for training, in contrast to AM and the full fine-tuning, which reduces considerable memory resources for solving each VRP (except the backbone Transformer for TSP). Besides transferring knowledge between problems with the same (Uniform) distribution, the cross-problem learning effectively can solve downstream VRPs with different distributions. Results on such cross-distribution applications are provided in Appendix B.
@@ 5.2 Ablation Study
Figure 4: Training curves with POMO.
@@ 5.3 Versatility
The cross-problem learning is versatile to different deep models for VRPs. We further deploy it with POMO Kwon et al. (2020), another Transformer based neural network which is developed from AM Kool et al. (2018) and achieves the state-of-the-art performance in solving TSP and CVRP. More details of POMO can be found in Appendix C. Specifically, we keep all the experimental settings with the same hyperparameters for training the original POMO. For our fine-tuning methods, we train POMO with the batch size 64, and the instances generated on the fly. For each epoch, we train 100,000 instances with them divided into batches. The deep model is trained with 100 epochs, when we observe that the training curves already converge. For the learning rate, we set $\eta=10^{-4}$ for the problem size $n=20$, and $\eta=10^{-5}$ for $n=50,100$ to avoid overfitting in the training process. As shown in Figure 4, the full fine-tuning generally converges faster and better than POMO trained from scratch. Furthermore, the inside tuning attains close performance to POMO, saving 78.87% parameters (i.e., 0.265M vs. 1.254M), which is more significant than the saving with AM (i.e., 71.33%).

### v3

@@ Abstract
Existing neural heuristics often train a deep architecture from scratch for each specific vehicle routing problem (VRP), ignoring the transferable knowledge across different VRP variants. This paper proposes the cross-problem learning to assist heuristics training for different downstream VRP variants. Particularly, we modularize neural architectures for complex VRPs into 1) the backbone Transformer for tackling the travelling salesman problem (TSP), and 2) the additional lightweight modules for processing problem-specific features in complex VRPs. Accordingly, we propose to pre-train the backbone Transformer for TSP, and then apply it in the process of fine-tuning the Transformer models for each target VRP variant. On the one hand, we fully fine-tune the trained backbone Transformer and problem-specific modules simultaneously. On the other hand, we only fine-tune small adapter networks along with the modules, keeping the backbone Transformer still. Extensive experiments on typical VRPs substantiate that 1) the full fine-tuning achieves significantly better performance than the one trained from scratch, and 2) the adapter-based fine-tuning also delivers comparable performance while being notably parameter-efficient. Furthermore, we empirically demonstrate the favorable effect of our method in terms of cross-distribution application and versatility. The code is available: https://github.com/Zhuoyi-Lin/Cross_problem_learning.
@@ 1 Introduction
  • As an early attempt, we propose cross-problem learning for solving VRPs. The pre-trained backbone Transformer for a basic VRP is used to foster the training of Transformers for downstream complex VRPs through fine-tuning.
@@
  • We develop fine-tuning methods for cross-problem learning, i.e., the full and adapter-based fine-tuning, by which we fine-tune the entire Transformer or only the lightweight problem-specific modules with small adapter networks.
  • Experiments show that the knowledge learned in the Transformer for TSP is well transferred to aid in training neural heuristics for other VRPs. While full fine-tuning gains better performance than Transformer trained from scratch, adapter-based fine-tuning methods attain comparable performance, with far fewer parameters trained and stored.
@@ 2.1 Neural Heuristics for VRP
Current neural heuristics for VRPs mostly learn policies to construct a solution in an autoregressive way. Among diverse neural construction heuristics Bello et al. (2017); Nazari et al. (2018); Dai et al. (2017); Ma et al. (2023), the first breakthrough is made by the Transformer Vaswani et al. (2017), which is cast as Attention Model (AM) to solve VRP variants such as TSP, OP and PCTSP Kool et al. (2018). After that, numerous neural heuristics are proposed based on AM for solving VRPs Xin et al. (2021a); Kwon et al. (2020); Hottung et al. (2022). Among them, Kwon et al. significantly ameliorate AM with policy optimization with multiple optima (POMO) Kwon et al. (2020), and enlighten considerable follow-up works with Transformers to solve VRPs Kim et al. (2022b); Kwon et al. (2021); Choo et al. (2022). As alternatives, some works endeavor to predict probabilities of edges in optimal solutions to VRPs, which are then leveraged to efficiently construct high-quality solutions with search algorithms such as greedy search, sampling, and Monte Carlo tree search Fu et al. (2021); Qiu et al. (2022); Hudson et al. (2022). On the other hand, neural improvement heuristics aim to learn policies to iteratively improve an initial but complete solution. Inspired by traditional local search algorithms, a few works learn local operations such as 2-opt and swap to promote the solution  Wu et al. (2021); Ma et al. (2021); Wang et al. (2021); Kim et al. (2022a). Inspired by specialized algorithms or solvers for VRPs, some literature attempts to learn repair or destroy operations in Large Neighborhood Search (LNS) Hottung and Tierney (2020); Gao et al. (2020). Deep learning is also applied to enhance solvers like Lin-Kernighan-Helsgaun (LKH) Xin et al. (2021b); Kim et al. (2021) and Hybrid Genetic Search (HGS) Santana et al. (2022).
@@ 2.2 Pre-Training, Fine-Tuning
The pre-training-then-fine-tuning paradigm has been widely verified effective in domains of Natural Language Processing (NLP) Dong et al. (2019); Devlin et al. (2018), Graph Neural Network (GNN) Qiu et al. (2020); Hu et al. (2020), Computer Vision (CV) Bao et al. (2022); Chen et al. (2020), etc. In this line of work, large models such as Transformers are first trained on basic tasks Dong et al. (2019); Ribeiro et al. (2020); Hu et al. (2020); You et al. (2020); Dosovitskiy et al. (2021); Touvron et al. (2021). Then the pre-trained model is fine-tuned to solve downstream tasks, with the neural network adapted by introducing additional structures or layers (which possess far fewer parameters than the pre-trained model) Houlsby et al. (2019); Pfeiffer et al. (2021); Zhao et al. (2020); Kitaev et al. (2020); Hu et al. (2021). In this manner, the pre-training can be conducted on considerable unlabeled data to learn general patterns and features, which are re-used in downstream tasks. Also, the knowledge learned during pre-training is transferred to fine-tuning process to facilitate more efficient training.
@@ 3.1 VRP Description
The solution to a VRP instance is feasible if it satisfies the problem-specific constraints. For example, a feasible solution to a TSP instance is a tour that goes through every node on the graph exactly once. A solution to OP is feasible if the length of a tour is not larger than a given value. Other VRPs like PCTSP and CVRP are defined on graphs in a similar fashion. Their details can be found in Appendix A. [^1] https://arxiv.org/abs/2404.11677
@@ 3.2 Transformer Based Construction Heuristics
$\nabla_{\theta}\mathcal{L}(\theta|\mathcal{G})=\mathbb{E}_{p_{\theta}(\pi|\mathcal{G})}[(\mu(\pi)-b(\mathcal{G}))\nabla_{\theta}\log p_{\theta}(\pi|\mathcal{G})],$ (1)
@@ 4.4 Adapter-Based Fine-Tuning for VRPs
The adapter network in the inside tuning comprises two linear layers with a LeakyReLU (LR) activation function in between such that:
$\text{APT}_{\text{in}}(h_{i})=\mathbf{W}_{1}^{\text{in}}\cdot\text{LR}(\mathbf{W}_{0}^{\text{in}}\cdot h_{i}+\mathbf{b}_{0}^{\text{in}})+\mathbf{b}_{1}^{\text{in}},$ (2)
@@
Instead of adjusting intermediate embeddings in the encoder, the side tuning adjusts the output from the encoder, with the following neural architecture:
$h^{{}^{\prime}}_{i}=\mathbf{BN}\left(h_{i}+\mathbf{MSL}_{i}\left(h_{0},\ldots,h_{n}\right)\right),$ (3)
$h_{i}^{{}^{\prime\prime}}=\mathbf{BN}\left(h^{{}^{\prime}}_{i}+\mathbf{FL}(h^{{}^{\prime}}_{i})\right),$ (4)
$\text{APT}_{\text{si}}(h_{i}^{{}^{\prime\prime}})=\mathbf{BN}(\text{LR}(\mathbf{W}_{1}^{\text{si}}\cdot\mathbf{BN}(\text{LR}(\mathbf{W}_{0}^{\text{si}}\cdot h_{i}^{{}^{\prime\prime}}+\mathbf{b}_{0}^{\text{si}}))+\mathbf{b}_{1}^{\text{si}})),$ (5)
@@
Inspired by Low-Rank Adaptation (LoRA) in Hu et al. (2021), we design a low-rank decomposition to adjust the output from any matrix in the pre-trained Transformer,
$\text{APT}_{\text{lo}}(h_{i})=\mathbf{W}_{p}\cdot h_{i}+\mathbf{B}_{p}\mathbf{A}_{p}\cdot h_{i},$ (6)
@@ 5 Experiments
Following the existing works, we generate the coordinates of nodes by uniformly sampling from the unit square $[0,1]\times[0,1]$. For OP, we set prizes all equal to 1. For PCTSP, we randomly sample prizes and penalties from Uniform distributions, following Kool et al. (2018). Meanwhile, we generate different-sized instances for OP and PCTSP, by setting the number of customer nodes to $n=$ 20, 50, and 100. We then use the Transformer based AM in Kool et al. (2018) to implement the cross-problem learning, and maintain most of the original experimental settings for a fair comparison. Specifically, we train the model with 2500 batches in each epoch and use 512 instances in each batch (except for the problem with $n=$ 100, where we use 256 instances). For each epoch, the training instances are generated on the fly, and we use 10,000 instances for the validation. Then the paired t-test ($\alpha=5\%$) is conducted after each epoch to replace the baseline model, according to the improvement significance. For all problems and sizes, we train 100 epochs for the full convergence of the training curves. The learning rate is set as $10^{-4}$ in all our methods, as the one for the original AM [^2] We tried to assign different learning rates to adapter networks and problem-specific modules for fine-tuning, which did not result in better performance.. The server with Intel Core i9-10940X CPU and NVIDIA GeForce RTX 2080Ti is used in our experiments.
We compare our method with AM Kool et al. (2018), the DRL based method with a strong performance on a variety of VRPs. We deploy the cross-problem learning based on its Transformer models. We also compare various baselines as follows. Regarding OP, we include 1) Gurobi, the state-of-the-art exact solver for various COPs Gurobi Optimization, LLC (2022); 2) Compass, the specialized Genetic Algorithm (GA) for OP Kobeaga et al. (2018); 3) Tsili, a classic randomized construction heuristic for OP with the manually crafted node probabilities  Tsiligirides (1984). As for PCTSP, besides Gurobi, we run two additional baselines: 4) ILS, the iterated local search metaheuristic that is widely used for routing problems; and 5) OR-Tools, a mature metaheuristics based solver for routing problems.
@@
    Bold means the best results among learning based methods. * means the results of the method reported in Kool et al. (2018).
Table 1: Comparison between cross-problem learning and baselines.
Refer to caption Refer to caption Refer to caption
Figure 3: Ablation study on fine-tuning methods and activation functions on PCTSP ($n=20,50,100$), respectively.
@@ 5.1 Comparison
We test all methods with 10,000 instances of OP and PCTSP, two typical VRPs that aim to maximize the total prize and minimize the distance plus the total penalty in the tour, respectively. We report the results of AM and Tsili by two decoding strategies, i.e., greedy search and sampling. For our method, we only report the better results derived by sampling. The sampling size is 1,280 for these methods, following Kool et al. (2018). For adapter-based fine-tuning, we only report results derived by inside tuning, since it significantly outstrips side tuning and LoRA (see Section 5.2). We compare the methods by different metrics, including the average objective value, average gap (to the best solutions), total runtime, and the number of trained parameters (for learning based methods). In Table 1, we can observe that our method with either full fine-tuning or inside tuning consistently outperforms AM with either greedy search or sampling, with significantly smaller gaps and similar runtime. While Gurobi achieves the optimal solutions for small problems, it cannot derive solutions for middle-sized OP and PCTSP ($n$=50, 100). Instead, we run Gurobi as a heuristic with a 30s limit, but it is inferior to our methods on OP ($n$=50, 100) or fails to obtain any solutions on PCTSP ($n$=100). While Compass and ILS are specialized algorithms for OP and PCTSP respectively, our method achieves comparable results in all cases, with most gap differences smaller than $1\%$. Moreover, our methods significantly surpass Tsili on OP ($n$=20, 50, 100), and outperform OR-Tools on PCTSP ($n$=100). Lastly, the inside tuning costs far fewer parameters for training, in contrast to AM and the full fine-tuning, which reduces considerable memory resources for solving each VRP (except the backbone Transformer for TSP). Besides transferring knowledge between problems with the same (Uniform) distribution, the cross-problem learning can effectively solve downstream VRPs with different distributions. Results on such cross-distribution applications are provided in Appendix B.
@@ 5.2 Ablation Study
Table 2: Inference results with POMO.
Refer to caption Refer to caption Refer to caption
Figure 4: Training curves with POMO on CVRP $(n=20,50,100)$, respectively.
@@ 5.3 Versatility
The cross-problem learning is versatile to different deep models for VRPs. We further deploy it with POMO Kwon et al. (2020), another Transformer based neural network which is developed from AM Kool et al. (2018) and achieves the state-of-the-art performance in solving TSP and CVRP. More details of POMO can be found in Appendix C. Specifically, we keep all the experimental settings with the same hyperparameters for training the original POMO. For our fine-tuning methods, we train POMO with the batch size 64, and the instances generated on the fly. For each epoch, we train 100,000 instances with them divided into batches. The deep model is trained with 100 epochs, when we observe that the training curves already converge. For the learning rate, we set $\eta=10^{-4}$ for the problem size $n=20$, and $\eta=10^{-5}$ for $n=50,100$ to avoid overfitting in the training process. As shown in Figure 4, the full fine-tuning generally converges faster and better than POMO trained from scratch. Furthermore, the inside tuning attains close performance to POMO, saving 79.59% parameters (i.e., 0.256M vs. 1.254M), which is more significant than the saving with AM (i.e., 71.33%).